OpenStudy (anonymous):

From a boat on the lake, the angle of elevation to the top of a cliff is 15°54'. If the base of the cliff is 967 feet from the boat, how high is the cliff (to the nearest foot)?

OpenStudy (loser66):

can you draw out the diagram?

OpenStudy (anonymous):

it doesnt have a diagram

OpenStudy (loser66):

YOU do it.

OpenStudy (anonymous):

yeah idk how to do it.. I suck at this!

OpenStudy (loser66):



OpenStudy (loser66):

and I am stuck !!! just help you until this. is there any thing pop up in your mind?

OpenStudy (anonymous):

I found this on yahoo arctan (sin (pi/2)) sin (pi/2) = 1 arctan 1 = pi/4 cliff's height H is the side of a right triangle having opposite angle 15°54' and the other side 967ft H = 967 tan(15°54') = 275ft does it look right?

OpenStudy (anonymous):

yayy thank you !! and how do I give you a medal? Its my first time on here

OpenStudy (loser66):

the concept is ok, but the calculation is not good H = 967*tan(15.54) = 268.89 near 269 ft, not 275ft I don't care about medal, friend

OpenStudy (anonymous):

15°54'not 15.54...

OpenStudy (anonymous):

thats what it is in the equation

OpenStudy (loser66):

hihihi, sure. I use calculator to calculate tan ( angle) not tan(number)

OpenStudy (anonymous):

wait i put the wrong solution from yahoo loll this is it my bad 54' is 54/60 of a degree to convert in decimals you must solve the proportion x/100 = 54/60 should be 90/100 = 0.9 so 15°54' = 15.9°